Three uniform spheres are located at the corners of an equilateral triangle. Each side of the triangle has a length of 1.39 m. Two of the spheres have a mass of 2.73 kg each. The third sphere (mass unknown) is released from rest. Considering only the gravitational forces that the spheres exert on each other, what is the magnitude of the initial acceleration of the third sphere?
Answer #1

F1 = GMm/r² {cos 60°, sin 60°}

F2 = GMm/r² {cos 120°, sin 120°}

R = F1 + F2

R = GMm/r² {cos 60°, sin 60°} + GMm/r² {cos 120°, sin 120°}

R = GMm/r² {0, √3}

acceleration on third sphere is

a = R/m

a = GMm/r² {0, √3}/m

a = 6.67* 10-11 * 2.73/1.392 {0, √3}

a = 1.63* 10-10 {0,1}

|a| =1.63* 10-10 m/s²